# Principal twistor models and asymptotic hyperkähler metrics

### 04 Mar 2026 v1

@@ 3.1. Definitions
After recalling the definition of a twistor space, we introduce the notions of a twistor model as a candidate for it, and a principal twistor model as a generalization of the latter. We then explain that a twistor model can be obtained from a principal twistor model by slicing along any real section.
@@
The following is a fundamental proposition on principal twistor models. We state that a twistor model can be obtained from a principal twistor model by slicing along any real section. First, we define a real section of the vector bundle $\mathcal{C}(2)$:
@@
The proof of the following proposition is straightforward and thus omitted.
@@ 3.2.1. Definition
The notation $M(1)$ implies that the gluing has weight $1$ with respect to the given $\mathbb{C}^{*}$-action.
@@ 3.2.3. Extension of quaternionic structure
$\left\{\begin{aligned} \hat{f}(x,u)&=\left(-\dfrac{1}{\bar{u}}\cdot f(x),-\dfrac{1}{\bar{u}}\right),\\ \hat{f}(x^{\prime},v)&=\left(-\dfrac{1}{\bar{v}}\cdot-f(x^{\prime}),-\dfrac{1}{\bar{v}}\right).\end{aligned}\right$. (3.4)
Here, f-f denotes the composition with the action of 1-1, i.e., $f\coloneqq-1\cdot f(x)$.
@@ 3.2.4. Extension of holomorphic differential forms
We show that when a holomorphic differential form on a manifold $M$ with a $\mathbb{C}^{*}$-action is compatible with the $\mathbb{C}^{*}$-action, a holomorphic differential form is naturally induced on the fiber bundle $M(1)$. The proof of the following proposition is straightforward and thus omitted.
@@ 3.2.5. Construction of the principal twistor model
Let $X$ be a conical symplectic variety with a good triple $(\lambda_{X},\omega_{X},j_{X})$. Assume that $X$ admits a crepant resolution $Y$. In this subsection, we show that a principal twistor model $\mathcal{Y}(1)$ can be naturally constructed by applying the $\mathbb{C}^{*}$-gluing construction to the triple induced on the universal Poisson deformation $\mathcal{Y}$ of the crepant resolution $Y$.
@@
Let $X$ be a conical symplectic variety with a good triple. Assume that $X$ admits a crepant resolution $Y$. The good triple on $X$ naturally induces the triple $(\tilde{\lambda}_{Y},\tilde{\omega}_{Y},\tilde{j}_{Y})$ on the total space of the universal Poisson deformation $\varphi_{Y}:\mathcal{Y}\rightarrow\mathcal{C}$ of the resolution $Y$ by Proposition 2.19. By applying the $\mathbb{C}^{*}$-gluing construction to this triple, we obtain a fiber bundle $\varphi_{Y}(1):\mathcal{Y}(1)\rightarrow\mathcal{C}(2)$, a real structure $\tau_{P}$, and a relative holomorphic symplectic $2$-form $\omega_{P}$ on $\mathcal{Y}(1)$. In this case, the tuple $(\varphi_{Y}(1):\mathcal{Y}(1)\rightarrow\mathcal{C}(2),\omega_{P},\tau_{P})$ is a principal twistor model.
@@
Figure 1. A conceptual diagram of the model $\mathcal{Y}(1)$, where the three coordinate axes correspond to $\mathcal{C}$, $Y$, and $\mathbb{P}^{1}$. The three faces correspond to $\mathcal{C}(2)$, $\mathcal{Y}$, and $Y(1)$.
Figure 2. A conceptual diagram showing the twistor model $Z_{s}$ associated with a real section $s$ embedded inside the model $\mathcal{Y}(1)$.
@@ 3.3. Universality of the Principal Twistor Model
Let $X$ be a conical symplectic variety. Assume that $X_{\mathrm{reg}}$ admits an algebraic hyperkähler cone metric $g_{0}$. Let $Y$ be a crepant resolution of $X$. In this section, we prove the universality of the principal twistor model, which states that the twistor space corresponding to any algebraic hyperkähler metric on the resolution $Y$ asymptotic to the metric $g_{0}$ is obtained by slicing the principal twistor model. To this end, we first show that the algebraic hyperkähler cone metric $g_{0}$ naturally induces the good triple on $X$. Next, we prove the aforementioned universality for the principal twistor model constructed from this good triple.
@@ 3.3.1. Twistor Cones and Hyperkähler Cone Metrics
We say that a twistor model $(Z\to\mathbb{P}^{1},\omega,\tau)$ is algebraic if it satisfies the following three conditions: Let $\mathbb{P}^{1}=\mathbb{C}^{+}\cup\mathbb{C}^{-}$. Let $Z^{\pm}\to\mathbb{C}^{\pm}$ denote the restrictions of the fiber bundle $Z$ to $\mathbb{C}^{\pm}$.
@@
    ALE gravitational instantons: These are typical examples of algebraic hyperkähler metrics (cf. Hitchin [4]).
@@
    Quiver varieties and toric hyperkähler manifolds: Hyperkähler metrics on Nakajima quiver varieties (cf. Nakajima [11]) and toric hyperkähler manifolds (cf. Bielawski and Dancer [1]) are also algebraic.
@@
Let $X$ be a conical symplectic variety. Assume that $X_{\mathrm{reg}}$ admits a metric $g_{0}$. We say that the metric $g_{0}$ is a cone metric on the regular locus $X_{\mathrm{reg}}$ if there exists a positive integer $k$ such that $\lambda_{r}^{*}g_{0}=r^{k}g_{0}$ for any action $\lambda_{r}$ with $r>0$.
@@
Let $X$ be a conical symplectic variety. Assume that $X_{\mathrm{reg}}$ admits an algebraic hyperkähler cone metric $g_{0}$, and let $(Z\to\mathbb{P}^{1},\omega,\tau,\{\ell_{x}\}_{x\in X_{\mathrm{reg}}})$ be the corresponding algebraic twistor space. Let $\omega_{0}$ be the restriction of the relative holomorphic symplectic form $\omega$ to the central fiber $Z_{0}\simeq X$. Note that the real structure $\tau$ restricts to an anti-holomorphic map $\tau|_{Z_{0}}:Z_{0}\to\overline{Z}_{\infty}$. Let $\tau_{0}$ be the anti-holomorphic involution on $X$ induced by this restriction via the natural identifications $Z_{\infty}\simeq X$. Denoting the $\mathbb{C}^{*}$-action on $X$ by $\lambda_{0}$, the tuple $(\lambda_{0},\omega_{0},\tau_{0})$ forms a good triple on $X$. Moreover, the twistor space $Z$ is isomorphic as a twistor model to the twistor cone $(X(1)\to\mathbb{P}^{1},\hat{\omega}_{0},\hat{\tau}_{0})$ constructed from this good triple (cf. Def. 3.25).
@@
Let $g_{0}$ be a hyperkähler cone metric on $X_{\mathrm{reg}}$ satisfying the assumptions of Proposition 3.32, and consider the pair $(\lambda_{0},\omega_{0})$. Let $\mathbb{P}^{1}=\mathbb{C}^{+}\cup\mathbb{C}^{-}$, and let $Z^{\pm}\to\mathbb{C}^{\pm}$ be the restrictions of the fiber bundle $Z$ to $\mathbb{C}^{\pm}$. Then, the Poisson deformation $Z^{\pm}$ on $\mathbb{C}^{\pm}$ is isomorphic to the trivial Poisson deformation $X\times\mathbb{C}$.
@@
    The anti-holomorphic maps $\tau^{+}$ and $\tau^{\prime}$$^{+}$ are conjugate: The two maps $\tau^{+}$ and $\tau^{\prime}$$^{+}$ are Poisson isomorphisms, coincide with the map $\tau_{0}$ on the central fiber, and coincide with the map $u\mapsto-\bar{u}$ on the base space $\mathbb{C}$. Moreover, they satisfy $(\tau^{+})^{4}=({\tau^{\prime}}^{+})^{4}=\mathrm{id}_{X\times\mathbb{C}}$. Consider the Poisson automorphisms $f,f^{\prime}$ on $X\times\mathbb{C}$ obtained by composing the two maps $\tau^{+},{\tau^{\prime}}^{+}$ with the complex conjugation on $X\times\mathbb{C}$. By Lemma 2.21 and Remark 2.22, the morphisms induced by the two morphisms $f,f^{\prime}$ on the formal Poisson deformation of $X\times\mathbb{C}$ are $\mathbb{C}^{*}$-equivariantly conjugate. Therefore, by considering their algebraization (cf. Cor. 2.25), it is shown that the two morphisms $f,f^{\prime}$ on $X\times\mathbb{C}$ are $\mathbb{C}^{*}$-equivariantly conjugate. From this, it follows that the two morphisms $\tau^{+}$ and $\tau^{\prime}$$^{+}$ are $\mathbb{C}^{*}$-equivariantly conjugate.
@@ 3.3.2. Universality of the Principal Twistor Model
Let $X$ be a conical symplectic variety. Assume that $X_{\mathrm{reg}}$ admits an algebraic hyperkähler cone metric $g_{0}$. Let $Y$ be a crepant resolution of $X$. In this subsection, we prove the universality of the principal twistor model, which states that the algebraic twistor space corresponding to any hyperkähler metric on the resolution $Y$ asymptotic to $g_{0}$ is obtained by slicing the principal twistor model.
@@
Let $X$ be a conical symplectic variety. Assume that $X_{\mathrm{reg}}$ admits a cone metric $g_{0}$. Let $Y$ be a crepant resolution of $X$. We say that a metric $g$ on the resolution $Y$ is asymptotic to the metric $g_{0}$ at infinity on $Y$ if it satisfies the following condition: Identifying $X_{\mathrm{reg}}$ with a subset of $Y$ naturally, fix an arbitrary point $x_{0}\in X_{\mathrm{reg}}$. Then, we have
@@
Here, $r\cdot x_{0}$ denotes the action of $r>0$ on $X$, and $||\cdot||_{g_{0}(r\cdot x_{0})}$ denotes the norm induced by the metric $g_{0}$ at the point $r\cdot x_{0}\in X_{\mathrm{reg}}$.
@@
The condition on the order of convergence of the limit is unnecessary.
@@
Let $X$ be a conical symplectic variety. Assume that $X_{\mathrm{reg}}$ admits an algebraic hyperkähler cone metric $g_{0}$. By Proposition 3.32, the metric $g_{0}$ induces the good triple on $X$. For this good triple, consider the principal twistor model $(\mathcal{Y}(1)\to\mathcal{C}(2),\omega_{P},\tau_{P})$ constructed in Proposition 3.23. Consider an algebraic hyperkähler metric $g$ on $Y$ that is asymptotic to the metric $g_{0}$ at infinity on $Y$. Then, there exists a unique real section $s$ of the vector bundle $\mathcal{C}(2)$ such that the twistor space $Z$ corresponding to the metric $g$ is isomorphic as a twistor model to the model $Z_{s}$ obtained by slicing the principal twistor model $\mathcal{Y}(1)$ along the real section $s$ by Proposition 3.9.
@@ 4.1.1. Structure of the Set of Twistor Lines
Let $X$ be a conical symplectic variety. Assume that $X_{\mathrm{reg}}$ admits an algebraic hyperkähler cone metric $g_{0}$. Let $Y$ be a crepant resolution of $X$. Consider the following setup and notation:
@@
Consider the setup in Notation 4.1. Then, the subset $\mathcal{T}=\sqcup_{s}\mathcal{T}_{s}\subset\Gamma(\mathbb{P}^{1},P)^{\tau_{P}}$ is a smooth manifold of real dimension $4n+3d$. Here, $4n=\dim_{\mathbb{R}}X$ and $d=\dim H^{2}(Y;\mathbb{C})$.
@@
Lemma 4.3 (P. Slodowy [18, p.49, §4.2 Remark]).
@@
    Consider the evaluation map $ev_{0}:\Gamma(\mathbb{P}^{1},P)^{\tau_{P}}\to\mathcal{Y}$ that maps a real section $\ell\in\Gamma(\mathbb{P}^{1},P)^{\tau_{P}}$ to its value $\ell(0)$. Fix a smooth trivialization $\mathcal{Y}\simeq Y\times\mathcal{C}$ of the universal Poisson deformation $\mathcal{Y}$ obtained by Lemma 4.3. Let $pr_{1}$ be the projection from the direct product $Y\times\mathcal{C}$ to $Y$. Then, we define a smooth map $ev:\Gamma(\mathbb{P}^{1},P)^{\tau_{P}}\to Y$ by $ev(\ell)=pr_{1}\circ\ell(0)$.
@@
The locally finite covering map $\Phi$ is not surjective in general.

### 12 Mar 2026 v2

@@ 3.1. Definitions
After recalling the definition of a twistor space, we introduce the notions of a twistor model as a candidate for it, and a principal twistor model as a generalization of the latter. We then explain that, given a principal twistor model, one can obtain a family of twistor models by slicing it along arbitrary real sections.
@@
Next, we define a real section of the vector bundle $\mathcal{C}(2)$:
@@
The following is a fundamental proposition on principal twistor models. We state that, for any choice of a real section $s$ of the vector bundle $\mathcal{C}(2)$, we obtain a twistor model $Z_{s}$ by slicing the principal twistor model along the section $s$. The proof of the following proposition is straightforward and thus omitted.
@@ 3.2.1. Definition
The notation $M(1)$ implies that the gluing has weight $1$ with respect to the given $\mathbb{C}^{*}$-action. Note also that the same construction applies equally well even if $M$ is a complex variety with singularities.
@@ 3.2.3. Extension of quaternionic structure
$\left\{\begin{aligned} \hat{f}(x,u)&=\left(-\dfrac{1}{\bar{u}}\cdot f(x),-\dfrac{1}{\bar{u}}\right),\\ \hat{f}(x^{\prime},v)&=\left(-\dfrac{1}{\bar{v}}\cdot-f(x^{\prime}),-\dfrac{1}{\bar{v}}\right),\end{aligned}\right$. (3.4)
where f-f denotes the composition with the action of 1-1, i.e., $f\coloneqq-1\cdot f(x)$.
@@ 3.2.4. Extension of holomorphic differential forms
We show that when a holomorphic differential form on a manifold $M$ with a $\mathbb{C}^{*}$-action is compatible with the $\mathbb{C}^{*}$-action, a relative holomorphic differential form is naturally induced on the fiber bundle $M(1)$. The proof of the following proposition is straightforward and thus omitted.
@@ 3.2.5. Construction of the principal twistor model
Let $X$ be a conical symplectic variety with a good triple $(\lambda,\omega,j)$. Assume that $X$ admits a crepant resolution $Y$. In this subsection, we show that a principal twistor model $\mathcal{Y}(1)$ can be naturally constructed by applying the $\mathbb{C}^{*}$-gluing construction to the triple induced on the universal Poisson deformation $\mathcal{Y}$ of the crepant resolution $Y$.
@@
Let $X$ be a conical symplectic variety with a good triple $(\lambda,\omega,j)$. Assume that $X$ admits a crepant resolution $Y$. The good triple on $X$ naturally induces the triple $(\tilde{\lambda}_{Y},\tilde{\omega}_{Y},\tilde{j}_{Y})$ on the total space of the universal Poisson deformation $\varphi_{Y}:\mathcal{Y}\rightarrow\mathcal{C}$ of the resolution $Y$ by Proposition 2.14. By applying the $\mathbb{C}^{*}$-gluing construction to this triple, we obtain a fiber bundle $\varphi_{Y}(1):\mathcal{Y}(1)\rightarrow\mathcal{C}(2)$, a real structure $\tau_{P}$, and a relative holomorphic symplectic $2$-form $\omega_{P}$ on $\mathcal{Y}(1)$. In this case, the tuple $(\varphi_{Y}(1):\mathcal{Y}(1)\rightarrow\mathcal{C}(2),\omega_{P},\tau_{P})$ is a principal twistor model.
@@
Figure 1. A conceptual diagram of the principal twistor model $\mathcal{Y}(1)$, where the three coordinate axes correspond to $\mathcal{C}$, $Y$, and $\mathbb{P}^{1}$. The three faces correspond to $\mathcal{C}(2)$, $\mathcal{Y}$, and $Y(1)$.
Figure 2. A conceptual diagram showing the twistor model $Z_{s}$ associated with a real section $s$ embedded inside the principal twistor model $\mathcal{Y}(1)$.
@@ 3.3. Universality of the Principal Twistor Model
Let $X$ be a conical symplectic variety. Assume that the regular locus $X_{\mathrm{reg}}$ admits an algebraic hyperkähler cone metric $g_{0}$. Let $Y$ be a crepant resolution of $X$. In this section, we prove the universality of the principal twistor model, which states that the twistor space corresponding to any algebraic hyperkähler metric on the resolution $Y$ asymptotic to the metric $g_{0}$ is obtained by slicing the principal twistor model. To this end, we first show that the algebraic hyperkähler cone metric $g_{0}$ naturally induces the good triple on $X$. Next, we prove the aforementioned universality for the principal twistor model constructed from this good triple.
@@ 3.3.1. Twistor Cones and Hyperkähler Cone Metrics
We say that a twistor model $(Z\to\mathbb{P}^{1},\omega,\tau)$ is algebraic if it satisfies the following three conditions: Suppose $\mathbb{P}^{1}=\mathbb{C}^{+}\cup\mathbb{C}^{-}$, and let $Z^{\pm}\to\mathbb{C}^{\pm}$ denote the restrictions of the fiber bundle $Z$ to $\mathbb{C}^{\pm}$.
@@
    ALE gravitational instantons: These are typical examples of algebraic hyperkähler metrics (cf. [4]).
@@
    Quiver varieties and toric hyperkähler manifolds: Hyperkähler metrics on Nakajima quiver varieties (cf. [9]) and toric hyperkähler manifolds (cf. [1]) are also algebraic.
@@
Let $X$ be a conical symplectic variety. Assume that the regular locus $X_{\mathrm{reg}}$ admits a metric $g_{0}$. We say that the metric $g_{0}$ is a cone metric on $X_{\mathrm{reg}}$ if there exists a positive integer $k$ such that $\lambda_{r}^{*}g_{0}=r^{k}g_{0}$ for any action $\lambda_{r}$ with $r>0$.
@@
Let $X$ be a conical symplectic variety. Assume that the regular locus $X_{\mathrm{reg}}$ admits an algebraic hyperkähler cone metric $g_{0}$, and let $(Z\to\mathbb{P}^{1},\omega,\tau,\{\ell_{x}\}_{x\in X_{\mathrm{reg}}})$ be the corresponding algebraic twistor space. Let $\omega_{0}$ be the restriction of the relative holomorphic symplectic form $\omega$ to the central fiber $Z_{0}\simeq X$. Note that the real structure $\tau$ restricts to an anti-holomorphic map $\tau|_{Z_{0}}:Z_{0}\to\overline{Z}_{\infty}$. Let $\tau_{0}$ be the anti-holomorphic involution on $X$ induced by this restriction via the natural identifications $Z_{\infty}\simeq X$. Denoting the $\mathbb{C}^{*}$-action on $X$ by $\lambda_{0}$, the tuple $(\lambda_{0},\omega_{0},\tau_{0})$ forms a good triple on $X$. Moreover, the twistor space $Z$ is isomorphic as a twistor model to the twistor cone $(X(1)\to\mathbb{P}^{1},\hat{\omega}_{0},\hat{\tau}_{0})$ constructed from this good triple (cf. Def. 3.25).
@@
Let $X$ be a conical symplectic variety. Assume that the regular locus $X_{\mathrm{reg}}$ admits an algebraic hyperkähler cone metric $g_{0}$, and let $(Z\to\mathbb{P}^{1},\omega,\tau,\{\ell_{x}\}_{x\in X_{\mathrm{reg}}})$ be the corresponding algebraic twistor space. Let $\mathbb{P}^{1}=\mathbb{C}^{+}\cup\mathbb{C}^{-}$, and let $Z^{\pm}\to\mathbb{C}^{\pm}$ be the restrictions of the fiber bundle $Z$ to $\mathbb{C}^{\pm}$. Then, the Poisson deformation $Z^{\pm}$ on $\mathbb{C}^{\pm}$ is isomorphic to the trivial Poisson deformation $X\times\mathbb{C}$.
@@
    The anti-holomorphic maps $\tau^{+}$ and $\tau^{\prime}$$^{+}$ are conjugate: The two maps $\tau^{+}$ and $\tau^{\prime}$$^{+}$ are Poisson isomorphisms, coincide with the map $\tau_{0}$ on the central fiber, and coincide with the map $u\mapsto-\bar{u}$ on the base space $\mathbb{C}$. Moreover, they satisfy $(\tau^{+})^{4}=({\tau^{\prime}}^{+})^{4}=\mathrm{id}_{X\times\mathbb{C}}$. Consider the Poisson automorphisms $f,f^{\prime}$ on $X\times\mathbb{C}$ obtained by composing the two maps $\tau^{+},{\tau^{\prime}}^{+}$ with the complex conjugation on $X\times\mathbb{C}$. By Lemma 2.16 and Remark 2.17, the morphisms induced by the two morphisms $f,f^{\prime}$ on the formal Poisson deformation of $X\times\mathbb{C}$ are $\mathbb{C}^{*}$-equivariantly conjugate. Therefore, by considering their algebraization (cf. Cor. 2.20), it is shown that the two morphisms $f,f^{\prime}$ on $X\times\mathbb{C}$ are $\mathbb{C}^{*}$-equivariantly conjugate. From this, it follows that the two morphisms $\tau^{+}$ and $\tau^{\prime}$$^{+}$ are $\mathbb{C}^{*}$-equivariantly conjugate.
@@ 3.3.2. Universality of the Principal Twistor Model
Let $X$ be a conical symplectic variety. Assume that the regular locus $X_{\mathrm{reg}}$ admits an algebraic hyperkähler cone metric $g_{0}$. Let $Y$ be a crepant resolution of $X$. In this subsection, we prove the universality of the principal twistor model, which states that the algebraic twistor space corresponding to any hyperkähler metric on the resolution $Y$ asymptotic to $g_{0}$ is obtained by slicing the principal twistor model.
@@
Let $X$ be a conical symplectic variety. Assume that $X_{\mathrm{reg}}$ admits a cone metric $g_{0}$. Let $Y$ be a crepant resolution of $X$. We say that a metric $g$ on the resolution $Y$ is asymptotic to the metric $g_{0}$ at infinity on $Y$ if it satisfies the following condition: Identifying the regular locus $X_{\mathrm{reg}}$ with a subset of $Y$ naturally, fix an arbitrary point $x_{0}\in X_{\mathrm{reg}}$. Then, we have
@@
where $r\cdot x_{0}$ denotes the action of $r>0$ on $X$, and $||\cdot||_{g_{0}(r\cdot x_{0})}$ denotes the norm induced by the metric $g_{0}$ at the point $r\cdot x_{0}\in X_{\mathrm{reg}}$.
@@
An assumption on the rate of convergence is not required.
@@
Let $X$ be a conical symplectic variety. Assume that the regular locus $X_{\mathrm{reg}}$ admits an algebraic hyperkähler cone metric $g_{0}$. By Proposition 3.32, the metric $g_{0}$ induces the good triple on $X$. For this good triple, consider the principal twistor model $(\mathcal{Y}(1)\to\mathcal{C}(2),\omega_{P},\tau_{P})$ constructed in Proposition 3.23. Consider an algebraic hyperkähler metric $g$ on $Y$ that is asymptotic to the metric $g_{0}$ at infinity on $Y$. Then, there exists a unique real section $s$ of the vector bundle $\mathcal{C}(2)$ such that the twistor space $Z$ corresponding to the metric $g$ is isomorphic as a twistor model to the model $Z_{s}$ obtained by slicing the principal twistor model $\mathcal{Y}(1)$ along the real section $s$ by Proposition 3.9.
@@ 4.1.1. Structure of the Set of Twistor Lines
Let $X$ be a conical symplectic variety. Assume that the regular locus $X_{\mathrm{reg}}$ admits an algebraic hyperkähler cone metric $g_{0}$. Let $Y$ be a crepant resolution of $X$. Consider the following setup and notation:
@@
Consider the setup in Notation 4.1. Then, the subset $\mathcal{T}=\sqcup_{s}\mathcal{T}_{s}\subset\Gamma(\mathbb{P}^{1},P)^{\tau_{P}}$ is a smooth manifold of real dimension $4n+3d$, where $4n=\dim_{\mathbb{R}}X$ and $d=\dim H^{2}(Y;\mathbb{C})$.
@@
Lemma 4.3 (P. Slodowy [15, p.49, §4.2 Remark]).
@@
    Consider the evaluation map $ev_{0}:\Gamma(\mathbb{P}^{1},P)^{\tau_{P}}\to\mathcal{Y}$ that maps a real section $\ell\in\Gamma(\mathbb{P}^{1},P)^{\tau_{P}}$ to its value $\ell(0)$. Fix a smooth trivialization $\mathcal{Y}\simeq Y\times\mathcal{C}$ of the universal Poisson deformation $\mathcal{Y}$ obtained by Lemma 4.3. Let $pr_{1}$ be the projection from the direct product $Y\times\mathcal{C}$ to $Y$. Then, we define a smooth map $ev:\Gamma(\mathbb{P}^{1},P)^{\tau_{P}}\to Y$ by $ev\coloneqq pr_{1}\circ ev_{0}$.
@@
The map $\Phi$ is not surjective in general.